A sample of a gas occupies 1640 mL at 70◦C and 740 torr. What volume would it occupy at the same pressure and 389◦C? Answer in units of mL.
1
Expert's answer
2019-12-09T08:04:22-0500

According to the ideal gas law:

P1×V1/T1 = P2×V2/T2,

where P1 and P2 - initial and final pressures, V1 and V2 - initial and final volumes, T1 and T2 - initial and final temperatures.

As the pressure is constant:

V1/T1 = V2/T2

From here,

V2 = V1×T2/T1

As V1 = 1640ml, T1 = 70◦C = 343 K, and T2 = 389◦C = 662 K:

V2 = 1640ml × 662 K / 343 K = 3165 ml


Answer: 3165 ml